# TARS [Uncaptioned image]: MinMax Token-Adaptive Preference Strategy for MLLM Hallucination Reduction

### v3

@@ Abstract
Multimodal large language models (MLLMs) enable vision-language reasoning, yet often generate plausible outputs that are factually incorrect or visually ungrounded, thereby compromising their reliability. Direct preference optimization (DPO) is a common strategy for correcting hallucinations by aligning model outputs with human preferences. However, existing DPO strategies typically treat hallucination-related preferences as fixed targets, relying on static and potentially biased supervision signals during training. This approach tends to overfit to superficial linguistic cues in preference data, leading to distributional rigidity and spurious correlations that impair grounding in causally relevant visual information. To overcome this limitation, we propose TARS, a token-adaptive preference strategy that reformulates DPO as a min–max optimization problem. TARS maximizes token-level distributional shifts under explicit semantic constraints to simulate alignment uncertainty, and simultaneously minimizes the expected preference loss under these controlled perturbations. This joint objective effectively preserves causal grounding while mitigating overfitting to preference patterns, thereby reducing hallucinations in multimodal reasoning. We evaluate TARS on multiple hallucination benchmarks and find consistently strong and robust performance. Using only 4.8k preference samples and no expert feedback, TARS reduces hallucination rates from 26.4% to 13.2% and decreases cognition value from 2.5 to 0.4, outperforming standard DPO and matching GPT-4o on several key metrics.
Refer to caption Refer to caption
Figure 1: Left: We present TARS, a token-adaptive preference strategy for mitigating hallucinations in MLLMs. TARS reformulates direct preference optimization (DPO) as a min-max objective that (1) minimizes behavioral misalignment via preference feedback and (2) maximizes adaptability through perturbations of visual-agnostic tokens. Right: Evaluation on LLaVA-v1.5-13B with preference optimization (PO) (Liu et al., 2023b) and various MLLMs under AMBER benchmark (Wang et al., 2023) shows that TARS surpasses PO baselines and matches GPT-4o (Hurst et al., 2024).
@@ 1 Introduction
Large language models (LLMs) demonstrate strong reasoning capabilities across a broad range of language tasks (Gandhi et al., 2023; Chen et al., 2023; Wang et al., 2025). Building on this foundation, multimodal large language models (MLLMs) integrate visual inputs to enable grounded understanding and vision-language reasoning (Tong et al., 2024; Huang et al., 2023; Driess et al., 2025). Although this integration broadens their applicability in various tasks (Jiang et al., 2024b; Shao et al., 2025), it also introduces key challenges, among which hallucinations are particularly prominent (Kim et al., 2024; Jiang et al., 2024a; Huang et al., 2024). Hallucinations in MLLMs refer to outputs that may appear plausible yet are factually incorrect or lack grounding in the visual context (Sarkar et al., 2025b; Gunjal et al., 2024). Addressing these failures is essential for improving the reliability, safety, and practical applicability of MLLMs in real-world applications.
@@
Figure 2: Motivation illustration for TARS. (a) and (b) illustrate standard DPO and our token-adaptive strategy. (c) shows a VQA example where DPO hallucinates, while TARS avoids ungrounded output. (d) and (e) visualize token-to-query attention maps during decoding. DPO over-attends to spurious tokens, while TARS attends to causally grounded visual-semantic cues.
Modern MLLMs are typically developed through a two-stage training pipeline that includes a knowledge-intensive pretraining phase (Dai et al., 2023; Bao et al., 2022; Zhang et al., 2024), followed by instruction tuning (Liu et al., 2024a; Wang et al., 2024b; Liu et al., 2023a). These stages endow the model with broad world knowledge and the ability to follow instructions in natural language (Bao et al., 2022; Liu et al., 2023b). Despite these capabilities, hallucinations often stem not from knowledge deficits, but from behavioral biases acquired during training that lead the model to generate plausible yet ungrounded outputs (Oh et al., 2024; Chen et al., 2025). To address these failures, preference optimization (PO) has become a prominent strategy for reducing hallucinations by aligning model outputs with human expectations (Schulman et al., 2017; Achiam et al., 2017). PO fine-tunes models using ranked response pairs derived from either human feedback (Sun et al., 2024; Ouyang et al., 2022) or AI-generated preferences (Yu et al., 2025; Sharma et al., 2024), providing direct supervision to reinforce grounded and faithful responses. Such methods have proven effective in mitigating hallucinations across diverse tasks.
Direct preference optimization (DPO) (Rafailov et al., 2023) has become a widely adopted method for hallucination reduction (Fu et al., 2025; Yang et al., 2025). While generally effective, current DPO methods rely heavily on preference data, which can cause models to overfit to shallow textual cues, such as high-frequency phrases or repetitive patterns in the training set (Huang et al., 2024; Liu et al., 2024b). Prior studies have observed that this overfitting leads MLLMs to generate responses that appear plausible but lack visual grounding, as illustrated in Figure 2(c). In our analysis (Figure 2(d)), we further find that DPO-trained models often assign high preference to outputs containing spurious correlation tokens, including prepositions or frequently mentioned objects, even when these elements are not visually grounded (Xie et al., 2024; Wang et al., 2024a). These observations reveal a core limitation of DPO: its reliance on static preference signals hinders generalization under shifting visual–textual contexts, leading to brittle cross-modal alignment and increased vulnerability to hallucination (Setlur et al., 2024; Fu et al., 2025). This rigidity also prevents the model from adapting to local semantic discrepancies, weakening its ability to ground responses in causally relevant visual information.
We formulate the challenge of distributional rigidity in preference optimization as a min-max token-adaptive alignment problem: maximizing distributional variation under semantic constraints, followed by minimizing the expected preference loss under these controlled perturbations. Specifically, we introduce perturbations to visual-agnostic tokens, textual elements with minimal cross-modal grounding, to simulate contextual variation and shift the input distribution without altering the semantic content. This setup enables the model to rely on causally grounded visual signals rather than superficial textual correlations, thereby mitigating hallucinations arising from overfitting to preference data (see Figure 2(b)). We refer to this approach as TARS (token-adaptive preference strategy), a lightweight and generalizable approach that enhances preference learning by introducing distribution-aware variability during fine-tuning.
To validate the effectiveness of our method, we evaluate TARS on LLaVA-v1.5 (Liu et al., 2023b) at 7B and 13B scales, comparing it against leading preference optimization approaches. Across a comprehensive suite of hallucination benchmarks spanning generative and discriminative tasks, TARS-enhanced LLaVA-v1.5 achieves consistently strong performance and matches GPT-4o (Hurst et al., 2024) in several settings. These results underscore the effectiveness of token-adaptive preference optimization in reducing hallucinations and advancing the trustworthiness of MLLMs.
Our contributions are as follows:
  • We reformulate preference learning as a min-max optimization objective that encourages token-level distributional shifts within semantic boundaries, while minimizing preference misalignment, thereby mitigating overfitting to rigid or spurious supervision signals.
  • We introduce TARS, a lightweight strategy that perturbs visual-agnostic tokens to simulate semantic variation, enhancing visual grounding and reducing hallucinations.
  • TARS achieves SOTA hallucination reduction using only 4.8k preference samples and no expert feedback, matching GPT-4o on several key metrics.
@@ 2 Preliminaries
Multimodal Large Language Models. MLLMs extend LLMs by incorporating visual inputs alongside textual prompts (Zhang et al., 2024). Formally, given an image $x$ and a prompt $q$, the model generates a textual response $y=(y_{1},\dots,y_{l})$ in an autoregressive manner (Liu et al., 2023b):
@@
where $\pi_{\theta}$ denotes the conditional generation policy parameterized by $\theta$. Given a textual input $q$ and a visual input $x$, the model tokenizes them into discrete sequences: textual tokens $q=\{q_{1},\dots,q_{m}\}$ and visual tokens $x=\{x_{1},\dots,x_{n}\}$. These tokens are mapped to embeddings and fused via cross-attention to integrate semantic signals from both modalities. The resulting context is then used by the decoder to autoregressively generate the output sequence (Dou et al., 2022; Yang et al., 2021).
Direct Preference Optimization. Direct preference optimization (DPO) (Rafailov et al., 2023) is an effective approach for aligning model behavior with human preferences. It bypasses explicit reward models by directly optimizing preferences from pairwise comparisons.
Traditional methods such as reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022) and AI feedback (RLAIF) (Yu et al., 2025) rely on training a scalar reward model $r_{\psi}(x,q,y)$ from preference pairs. This reward model is typically trained using the Bradley-Terry formulation (Bradley & Terry, 1952):
@@
where $(x,q,y_{w},y_{r})$ is sampled from the preference data distribution $\mathcal{D}$, $\sigma(z)=\frac{1}{1+\exp(-z)}$ denotes the sigmoid function. $y_{w}$ and $y_{r}$ denote the preferred and dispreferred responses, respectively. $r_{\psi}(x,q,y)$ is trained to maximize the log-likelihood of correctly ranking the preferred response:
$\underset{r_{\psi}}{\min}\,\mathbb{E}_{(x,q,y_{w},y_{r})\sim\mathcal{D}}\left[-\log\sigma\left(r_{\psi}(x,q,y_{w})-r_{\psi}(x,q,y_{r})\right)\right],$ (3)
@@
$\underset{\pi_{\theta}}{\min}\,\mathbb{E}_{(x,q)\sim\mathcal{D},\,y^{*}\sim\pi_{\theta}(y^{*}\mid x,q)}\Big{[}-\Big{(}r_{\psi}(x,q,y^{*})-\alpha\cdot\mathbb{D}_{\mathrm{KL}}\left(\pi_{\theta}(y^{*}\mid x,q)\,\middle\|\,\pi_{\text{ref}}(y^{*}\mid x,q)\right)\Big{)}\Big{]},$ (4)
where $\alpha$ controls the strength of regularization, which ensures alignment with the learned preferences. Rather than relying on the explicitly trained reward model, DPO (Rafailov et al., 2023) simplifies the learning process by leveraging the insight that the optimal policy can be expressed in closed form using relative log-likelihoods under $\pi_{\theta}$ and $\pi_{\text{ref}}$:
$\underset{\pi_{\theta}}{\min}\,\mathbb{E}_{(x,q,y_{w},y_{r})\sim\mathcal{D}}\Big{[}-\log\sigma\Big{(}\alpha\log\frac{\pi_{\theta}(y_{w}\mid x,q)}{\pi_{\text{ref}}(y_{w}\mid x,q)}-\alpha\log\frac{\pi_{\theta}(y_{r}\mid x,q)}{\pi_{\text{ref}}(y_{r}\mid x,q)}\Big{)}\Big{]},$ (5)
This formulation enables direct policy optimization from preference pairs, aligning the output probabilities of MLLMs with human preferences.
@@ 3 Method
We propose a token-adaptive min-max strategy with perturbations on visual-agnostic tokens and a frequency-based regularizer for improved alignment (See Figure 3).
@@ 3.1 Min-Max Reformulation of DPO
To address the limitations of traditional DPO, we reformulate preference optimization as a token-adaptive min-max game. The inner maximization introduces controlled token-level perturbations $\varphi(\cdot)$ to induce input distribution shifts, while the outer minimization aligns the policy $\pi_{\theta}$ with preference signals. Formally, we define the min–max preference objective as:
@@
where $\varphi$ is a token-level perturbation function constrained to visually agnostic tokens, i.e., $\Phi(\mathcal{A})=\{\varphi\mid\{i\mid\varphi(q_{i})\neq q_{i}\}\subseteq\mathcal{A}(x,q)\}$. This min–max objective promotes preference alignment under distributional shifts, helping to mitigate spurious correlations and reduce hallucinated outputs.
@@ 3.2 Maximizing with Token Perturbations
As shown in Equation 5, DPO aligns models with preferred responses via log-likelihood ratios against a reference model. However, we observe that this formulation can encourage overfitting to superficial patterns such as frequent phrases, stylistic tokens, which we find reduce effective alignment with the visual context in multimodal settings (Setlur et al., 2024; Fu et al., 2025).
@@
where $\varphi(q_{i})$ is constructed using masking or synonym substitution. This approximation simulates worst-case alignment uncertainty while preserving semantic integrity.
To preserve semantics, we restrict changes to visual-agnostic tokens with minimal impact on cross-modal alignment. We compute token-level visual relevance as the similarity between visual features $\mathcal{G}_{v}(x)$ and token embeddings $\mathcal{G}_{t}(q_{i})$. We then identify a set $\mathcal{A}$ of $N_{t}$ visually agnostic tokens with the lowest cross-modal alignment scores:
$\mathcal{A}=\text{Top}_{N_{t}}\left(-\mathcal{G}_{v}(x)\mathcal{G}_{t}(q_{i})^{T}\right),N_{t}=\left\lfloor\omega\cdot\Delta P^{-1}\right\rfloor+1,$ (9)
where $\lfloor\cdot\rfloor$ denotes the floor operation, $\omega$ is a scaling coefficient. $P=-\mathcal{G}_{v}(x)\mathcal{G}_{t}(q_{i})^{T}$ is the similarity score matrix, $\Delta P=\max_{j}P_{j}-\max_{k\neq j}P_{k}$ quantifies the predictive uncertainty of MLLMs. We adapt $N_{t}$ inversely to this margin: confident predictions lead to fewer perturbations, while greater uncertainty induces broader variation.

### v4

# 4AWS AI Lab, Amazon  5Hupan Laboratory

@@ Abstract
Multimodal large language models (MLLMs) are prone to hallucinations, generating plausible but visually ungrounded outputs, partly because direct preference optimization (DPO) overfits to superficial linguistic cues under static preference supervision. We propose TARS, a token-adaptive preference strategy that reformulates DPO as a principled min-max optimization problem. The inner maximization selectively perturbs visual-agnostic tokens to induce worst-case distributional shifts, while the outer minimization enforces alignment with causal visual signals rather than surface-level patterns. A novel spectral alignment loss further regularizes hidden representations in the frequency domain via the Fast Fourier Transform (FFT), preserving global semantic structure without rigid token-level correspondence. We evaluate TARS across multiple hallucination benchmarks. Using only 4.8k preference samples without expert feedback, TARS reduces hallucination rates from 26.4% to 13.2% and cognition scores from 2.5 to 0.4, outperforming standard DPO by a large margin. Notably, TARS surpasses $5\times$ LLM-based data augmentation trained on 28.8k samples (Hal-Rate: 16.0% vs. 13.2%), demonstrating that reshaping the optimization landscape via adversarial token perturbation is fundamentally more effective than scaling training data. TARS further narrows the gap with GPT-4o on key metrics.
[Uncaptioned image] [Uncaptioned image]
Figure 1: Left: We present TARS, a token-adaptive preference strategy for mitigating hallucinations in MLLMs. TARS reformulates direct preference optimization (DPO) as a principled min-max optimization objective: (1) minimizes behavioral misalignment via structured preference feedback supervision and (2) maximizes distributional adaptability through controlled perturbations of visual-agnostic tokens. Right: Comprehensive evaluation on LLaVA-v1.5-13B with preference optimization (PO) [llava_origin] and various state-of-the-art MLLMs under the AMBER benchmark [wang2023amber] shows that TARS consistently surpasses PO baselines, yielding results competitive with GPT-4o [hurst2024gpt].
@@ 1 Introduction
Multimodal large language models (MLLMs) extend the reasoning capabilities of LLMs [gandhi2023understanding, chen2023large, wang2025reasoning] to visual inputs, enabling grounded vision-language understanding [tong2024cambrian, huang2023language, feng2025can]. Despite strong performance across diverse tasks [jiang2024marvel, shao2025holitom], MLLMs remain prone to hallucinations, producing outputs that appear plausible but are factually incorrect or lack visual grounding [kim2024code, jiang2024hallucination, huang2024opera, sarkarmitigating, gunjal2024detecting]. Mitigating such failures is crucial for deploying reliable MLLMs.
Modern MLLMs follow a two-stage pipeline of knowledge pretraining [instruct_blip, bao2022vlmo, 10445007] and instruction tuning [liu2024improved, liu2023aligning]. Hallucinations often stem not from knowledge deficits but from behavioral biases that produce plausible yet ungrounded outputs [oh2024erbench, chen2025perturbollava]. Preference optimization (PO) addresses this by fine-tuning with ranked response pairs from human [RLHF, RLHF_1] or AI feedback [RLAIF, sharma2024critical], aligning outputs with factual expectations [PPO, achiam2017constrained]. Direct preference optimization (DPO) [rafailov2023direct] is widely used for hallucination reduction [CHIP, OPA_DPO], yet current methods can overfit to shallow textual cues such as high-frequency phrases [huang2024opera, liu2024paying], generating plausible but visually ungrounded responses (Fig.˜2(c)). Our analysis (Fig.˜2(d)) further reveals that DPO-trained models assign high preference to outputs with spurious correlation tokens (e.g., prepositions or frequently mentioned objects) that lack visual grounding [xie2024v, wang2024mdpo]. This reliance on static preference signals hinders generalization under shifting visual-textual contexts, leading to brittle alignment and increased hallucination [setlur2024rl, CHIP].
@@
Figure 2: Motivation illustration for TARS. (a) and (b) illustrate standard DPO and our token-adaptive perturbation strategy. (c) shows a VQA example where DPO hallucinates, while TARS effectively avoids ungrounded output. (d) and (e) visualize token-to-query attention maps during autoregressive decoding. DPO over-attends to spurious tokens, while TARS attends to causally grounded visual-semantic cues.
We formulate this challenge as a min-max token-adaptive alignment problem: maximizing distributional variation under semantic constraints, then minimizing preference loss under these perturbations. A natural alternative is data augmentation, which diversifies the preference distribution by generating additional training pairs. However, our experiments (Sec.˜4.7) reveal that even $5\times$ LLM-based augmentation (28.8k samples) underperforms TARS using only the original 4.8k samples (Hal-Rate: 16.0% vs. 13.2%), demonstrating that reshaping the optimization landscape via adversarial token perturbation is fundamentally more effective than merely scaling training data. The key distinction is that data augmentation merely enriches the data distribution, whereas TARS actively reshapes the optimization landscape: the inner maximization dynamically generates worst-case token configurations per sample, forcing the outer minimization to learn alignment features that are invariant to distributional shifts rather than memorizing surface-level patterns.
Specifically, we perturb visual-agnostic tokens, i.e., textual elements with minimal cross-modal grounding, to shift the input distribution without altering semantic content. This forces the model to rely on causally grounded visual signals rather than superficial correlations (Fig.˜2(b)). Supervising perturbed representations requires maintaining semantic consistency without rigid token-level correspondence. Standard spatial constraints (e.g., $\ell_{2}$ or contrastive losses) indiscriminately penalize any positional deviation equally, reintroducing the spurious correlations our perturbation aims to eliminate [zhou2024explore, tianblack]. We instead propose a novel spectral alignment loss via the Fast Fourier Transform (FFT). Under the shift property of the Discrete Fourier Transform, a local token perturbation at position $l$ contributes a bounded additive term proportional to $e^{-2\pi ikl/L}$ to each frequency bin $k$, regardless of $l$ [cooley1965algorithm]. This means spectral magnitudes vary smoothly under local perturbation, whereas spatial metrics ($\ell_{2}$, cosine) exhibit sharp discontinuities at the perturbed positions. Meanwhile, the dominant low-frequency components capture global semantic structure (e.g., topic, intent, visual grounding), while high-frequency components encode position-dependent lexical details. This makes frequency-domain alignment a theoretically principled choice for preserving global semantic invariance while tolerating the local distributional shifts inherent in our min-max formulation (Sec.˜3.3). [^1] Empirical comparisons between spectral and spatial alignment strategies are provided in the Appendix, Section 9.4.
We refer to this approach as TARS (token-adaptive preference strategy), a lightweight and generalizable approach that enhances preference learning by combining adaptive token perturbation with frequency-domain spectral regularization. We evaluate TARS on LLaVA-v1.5 [llava_origin] at 7B and 13B scales across generative and discriminative hallucination benchmarks. TARS achieves consistently strong performance across all benchmarks and matches GPT-4o [hurst2024gpt] in several settings, underscoring the effectiveness of token-adaptive preference optimization. Our contributions are as follows:
  • We reformulate preference learning as a min-max optimization that dynamically generates worst-case token perturbations per sample, reshaping the optimization landscape rather than augmenting data. With only 4.8k samples, TARS outperforms $5\times$ LLM-based augmentation (28.8k samples).
  • We propose a spectral alignment loss via FFT that preserves global semantic structure in dominant low-frequency components while tolerating local distributional shifts from token perturbation, avoiding spurious positional correlations reintroduced by spatial constraints (e.g., $\ell_{2}$, cosine).
  • We present TARS, combining adaptive token perturbation with spectral regularization. Using only 4.8k samples without expert feedback, TARS achieves state-of-the-art hallucination reduction, matching or surpassing GPT-4o on key metrics while preserving general reasoning capabilities.
@@ 2 Preliminaries
Multimodal Large Language Models. MLLMs extend LLMs by incorporating visual inputs alongside textual prompts [10445007]. Formally, given an image $x$ and a prompt $q$, the model generates a textual response $y=(y_{1},\dots,y_{l})$ in an autoregressive manner [llava_origin]:
@@
where $\pi_{\theta}$ denotes the conditional generation policy parameterized by $\theta$. Given a textual input $q$ and a visual input $x$, the model tokenizes them into discrete sequences: textual tokens $q=\{q_{1},\dots,q_{m}\}$ and visual tokens $x=\{x_{1},\dots,x_{n}\}$. In practice, a pretrained vision encoder extracts patch-level features from the image, which are projected into the language model’s embedding space through a learnable alignment module. These tokens are mapped to embeddings and fused via cross-attention to integrate semantic signals from both modalities. The resulting multimodal context is then used by the decoder to autoregressively generate the output sequence [dou2022empirical, yang2021causal].
Direct Preference Optimization. Direct preference optimization (DPO) [rafailov2023direct] is an effective and widely adopted approach for aligning model behavior with human preferences. It bypasses explicit reward models by directly optimizing preferences from pairwise comparisons.
Traditional methods such as reinforcement learning with human feedback (RLHF) [RLHF_1] and AI feedback (RLAIF) [RLAIF] rely on training a scalar reward model $r_{\psi}(x,q,y)$ from preference pairs. This reward model is typically trained using the Bradley-Terry formulation [bradley1952rank]:
@@
where $(x,q,y_{w},y_{r})$ is sampled from the preference data distribution $\mathcal{D}$, and $\sigma(z)=\frac{1}{1+\exp(-z)}$ denotes the sigmoid function. $y_{w}$ and $y_{r}$ denote the preferred and dispreferred responses, respectively. The reward model $r_{\psi}(x,q,y)$ is then trained to maximize the log-likelihood of correctly ranking the preferred response over the dispreferred one during optimization:
$\underset{r_{\psi}}{\min}\,\mathbb{E}_{(x,q,y_{w},y_{r})\sim\mathcal{D}}\left[-\log\sigma\left(r_{\psi}(x,q,y_{w})-r_{\psi}(x,q,y_{r})\right)\right].$ (3)
@@
$\underset{\pi_{\theta}}{\min}\;\mathbb{E}_{(x,q)\sim\mathcal{D},\,y^{*}\sim\pi_{\theta}}\big[-\big(r_{\psi}(x,q,y^{*})-\alpha\cdot\mathbb{D}_{\mathrm{KL}}(\pi_{\theta}(y^{*}\mid x,q)\|\pi_{\text{ref}}(y^{*}\mid x,q))\big)\big].$ (4)
where $\alpha$ controls the strength of KL regularization, which ensures alignment with the learned preferences. Rather than relying on the explicitly trained reward model, DPO [rafailov2023direct] simplifies the learning process by leveraging the insight that the optimal policy can be expressed in closed form using relative log-likelihoods under $\pi_{\theta}$ and $\pi_{\text{ref}}$:
$\underset{\pi_{\theta}}{\min}\;\mathbb{E}_{(x,q,y_{w},y_{r})\sim\mathcal{D}}\big[-\log\sigma\big(\alpha\log\frac{\pi_{\theta}(y_{w}\mid x,q)}{\pi_{\text{ref}}(y_{w}\mid x,q)}-\alpha\log\frac{\pi_{\theta}(y_{r}\mid x,q)}{\pi_{\text{ref}}(y_{r}\mid x,q)}\big)\big].$ (5)
This formulation enables direct policy optimization from preference pairs, aligning the output probabilities with human preferences and improving alignment stability and sample efficiency compared to RL-based approaches.
@@ 3 Method
We propose a token-adaptive min-max strategy with perturbations on visual-agnostic tokens and a frequency-based regularizer for improved alignment. An overview is shown in Fig.˜3, and the detailed algorithm is provided in the Appendix. [^2] Full pseudocode is given in Appendix, Section 8. We first define notation used throughout: $x$ denotes the visual input, $q=\{q_{1},\dots,q_{m}\}$ the tokenized textual prompt, $y_{w}$ and $y_{r}$ the preferred and dispreferred responses, $\pi_{\theta}$ the trainable policy, $\pi_{\text{ref}}$ the frozen reference policy, and $\mathcal{D}$ the preference dataset.
@@ 3.1 Min-Max Reformulation of DPO
To address the limitations of traditional DPO, we reformulate preference optimization as a token-adaptive min-max game. The inner maximization introduces a controlled token-level perturbation function $\varphi(\cdot)$, which modifies selected tokens in the prompt $q$ to induce input distribution shifts. The outer minimization then aligns the policy $\pi_{\theta}$ with preference signals under these perturbations. Formally, we define the min–max preference objective as:
@@
where $\varphi$ is constrained to modify only visually agnostic tokens. Specifically, $\Phi(\mathcal{A})$ denotes the set of admissible perturbation functions: $\Phi(\mathcal{A})=\{\varphi\mid\{i\mid\varphi(q_{i})\neq q_{i}\}\subseteq\mathcal{A}(x,q)\}$, where $\mathcal{A}(x,q)$ is the set of token indices identified as visually agnostic (defined in Eq.˜9). This min–max objective promotes preference alignment under distributional shifts, helping to mitigate spurious correlations and reduce hallucinated outputs.
@@ 3.2 Maximizing with Token Perturbations
As shown in Eq.˜5, DPO aligns models with preferred responses via log-likelihood ratios against a reference model. However, we observe that this formulation can encourage overfitting to superficial patterns such as frequent phrases and stylistic tokens, which in turn reduce effective alignment with the visual context and hinder robust multimodal understanding [setlur2024rl, CHIP].
@@
where $\varphi(q_{i})$ denotes the perturbed token, constructed using either masking (replacing with [MASK]) or synonym substitution (implementation details are in Appendix, Section 7). $\mathbb{I}[\cdot]$ is the indicator function. This approximation simulates worst-case alignment uncertainty while preserving semantic integrity.
To preserve semantics, we restrict changes to visual-agnostic tokens with minimal impact on cross-modal alignment. We compute token-level visual relevance using a visual encoder $\mathcal{G}_{v}(\cdot)$ and a text encoder $\mathcal{G}_{t}(\cdot)$ (both instantiated by the CLIP encoder [radford2021learning]) as the cosine similarity between visual features $\mathcal{G}_{v}(x)\in\mathbb{R}^{d}$ and each token embedding $\mathcal{G}_{t}(q_{i})\in\mathbb{R}^{d}$. We then identify a set $\mathcal{A}$ of $N_{t}$ visually agnostic tokens with the lowest cross-modal alignment scores:
$\mathcal{A}=\text{Top}_{N_{t}}\!\big(-\,\mathcal{G}_{v}(x)\mathcal{G}_{t}(q_{i})^{\top}\big),\quad N_{t}=\lfloor\omega\cdot\Delta P^{-1}\rfloor+1.$ (9)
where $\lfloor\cdot\rfloor$ denotes the floor operation and $\omega$ is a scaling coefficient controlling perturbation intensity. The matrix $P\in\mathbb{R}^{m}$ is the negated similarity score vector, with $P_{i}=-\mathcal{G}_{v}(x)\cdot\mathcal{G}_{t}(q_{i})^{T}$. The confidence margin $\Delta P=\max_{j}P_{j}-\max_{k\neq j}P_{k}$ quantifies the predictive uncertainty of the cross-modal alignment: confident predictions (large $\Delta P$) lead to fewer perturbations, while greater uncertainty (small $\Delta P$) induces broader variation. An empirical comparison of this adaptive selection strategy against random and uniform perturbation baselines is provided in the Appendix, Section 9.5.